OpenStudy (anonymous):

5 square root of 40?

OpenStudy (anonymous):

This simply means 5 multiplied by whatever the value of the squared root of 40 is.

OpenStudy (anonymous):

.. 5(6.32)= approx 31.6

OpenStudy (anonymous):

is that what you wanted?

OpenStudy (anonymous):

Or perhaps this: \[5\sqrt{40}=5\sqrt{5*2*2*2}=5\left(2 \sqrt{10}\right)=10 \sqrt{10} \]